A charged oil drop of radius 1.3 × 10-6m is prevented from falling under gravity by the vertical electric field between two horizontal parallel plates charged to a difference of potential of 8340 V. The distance between the plates is 16 mm, and the density of oil is 920 kgm-3. Calculate the magnitude of the charge on the drop


1
Expert's answer
2021-01-10T18:31:31-0500

There are two forces that act on the charged oil drop: the electric force "F_e" directed upward and the force of gravity "F_g" directed downward. In order to oil drop to be prevented from falling, the force of gravity must be balanced by the electric force:


"F_e=F_g,""qE=mg."

From this equation, we can find the magnitude of the charge on the drop:


"q=\dfrac{mg}{E}. (1)"

We can find the electric field, "E", from the formula:


"E=\dfrac{V}{d}, (2)"

here, "V=8340\ V" is the potential difference between two horizontal parallel plates and "d=1.6\cdot10^{-2}\ m" is the distance between the plates.

We can find the mass of the oil drop from the definition of the density:


"\rho=\dfrac{m}{V},""m=\rho V,"

here, "\rho=920\ \dfrac{kg}{m^3}" is the density of oil, "V" is the volume of the oil drop.

The volume of the oil drop can be found from the formula:


"V=\dfrac{4}{3}\pi r^3,"

here, "r=1.3\cdot10^{-6}\ m" is the radius of the oil drop.

Then, for the mass of the oil drop we have:


"m=\rho V=\dfrac{4}{3}\pi\rho r^3. (3)"

Finally, substituting equations (2) and (3) into equation (1), we get:


"q=\dfrac{\dfrac{4}{3}\pi\rho r^3gd}{V},"

"q=\dfrac{\dfrac{4}{3}\cdot\pi\cdot920\ \dfrac{kg}{m^3}\cdot(1.3\cdot10^{-6}\ m)^3\cdot 9.8\ \dfrac{m}{s^2}\cdot 1.6\cdot10^{-2}\ m}{8340\ V}=1.6\cdot10^{-19}\ C."

Answer:

"q=1.6\cdot10^{-19}\ C."